# On the Fragility of Multimodal Perception to Temporal Misalignment in Autonomous Driving

### v1

@@ Abstract
Multimodal fusion (MMF) plays a critical role in the perception of autonomous driving, which primarily fuses camera and LiDAR streams for a comprehensive and efficient scene understanding. However, its strict reliance on precise temporal synchronization exposes it to new vulnerabilities. In this paper, we introduce DejaVu, a novel attack that exploits network-induced delays to create subtle temporal misalignments across sensor streams, severely degrading downstream MMF-based perception tasks. Our comprehensive attack analysis across different models and datasets reveals these sensors’ task-specific imbalanced sensitivities: object detection is overly dependent on LiDAR inputs, while object tracking is highly reliant on the camera inputs. Consequently, with a single-frame LiDAR delay, an attacker can reduce the car detection mAP by up to 88.5%, while with a three-frame camera delay, multiple object tracking accuracy (MOTA) for car drops by 73%. To detect such attacks, we propose Aion, a defense patch that can work alongside the existing perception model to monitor temporal alignment through cross-modal temporal consistency. Aion leverages multimodal shared representation learning and dynamic time warping to determine the path of temporal alignment and calculate anomaly scores based on the alignment. Our thorough evaluation of Aion shows it achieves AUROC scores of 0.92–0.98 with low false positives across datasets and model architectures, demonstrating it as a robust and generalized defense against the temporal misalignment attacks.
@@ I Introduction
Autonomous driving (AD) is designed to navigate and interact with complex environments—a capability fundamentally reliant on a comprehensive understanding of its surroundings. The integration of heterogeneous sensors, such as cameras, LiDARs, and radar, that capture data from different modalities allows an accurate perception with enhanced accuracy [1]. Each individual modality has unique strengths; for example, cameras capture rich semantic details, LiDAR provides accurate depth measurements, and radar particularly excels in detecting speed, even in adverse weather conditions [2]. However, these sensors also face inherent limitations, such as a camera’s sensitivity to lighting variations, LiDAR’s lack of texture information, and radar’s sparsity, which can compromise performance when used independently [3, 4, 5]. On the other hand, multimodal fusion (MMF) refers to the process of integrating such heterogeneous unimodal sensor data to generate a single and comprehensive representation of the surroundings for a more accurate perception [6], ensuring efficient downstream tasks such as object detection, tracking, and segmentation. MMF not only compensates for individual sensor weaknesses but also enhances the robustness of the overall system—a critical assumption that underpins its effectiveness in AD applications [7].
Although MMF has demonstrated considerable effectiveness, it remains a complex problem and a fundamental research challenge due to the intrinsic heterogeneity of sensing modalities, due to unique data formats, resolutions, and noise characteristics [8, 9]. This challenge is further compounded by the critical requirement of temporal alignment during sensor fusion, which is complicated by varying sampling rates across sensors, stemming from underlying hardware limitations [10]. In safety-critical cyber-physical systems like AD, it is obligatory for real-time perception, decision-making, and control. On the other hand, temporal misalignments in sensor data acquisition—such as unexpected delays—can lead to significant degradation in the perception tasks, manifesting as object misdetection, inaccurate localization, and hence, scene misinterpretation. These perceptual inaccuracies undermine the reliability and safety of AD, adversely affecting core functionalities such as longitudinal and lateral control, maneuver execution, safety-critical interventions, etc. [11].
These AD functions demand a perfectly synchronized sensor network with low latency to make timely control decisions [12]. In that regard, Automotive Ethernet (AE) has emerged as a key networking technology for in-vehicular networks, offering data rates from 10 Mbps to 10 Gbps [13]. Moreover, equipped with Time-Sensitive Networking (TSN), which uses generalized Precision Time Protocol (gPTP) for precise clock synchronization, ensures deterministic and real-time communication with bounded latency, which is essential for sensor fusion and control in AD applications [14, 15].
@@
Figure 1: Impact of DejaVu attack on 3D object detection. (a-b) show two benign scenarios without any temporal misalignment, and hence, accurate object detection. (c-d) Illustrate different temporal alignments between camera and LiDAR inputs and highlight how delayed sensor data can lead to incorrect detections, either by detecting non-existent objects (false positives) or missing present ones (false negatives). In (c), the MMF prioritizes the (delayed) LiDAR data and predicts three objects, including a pedestrian who is not present in the current camera view, resulting in a false detection from the camera’s perspective. However, in (d), the MMF still prioritizes the (updated) LiDAR data and predicts two objects, excluding the pedestrian who is present in the current camera view, resulting in a missed detection from the camera’s perspective. In both temporal misalignment attack cases, MMF biased its fusion toward LiDAR, failing to account for semantic discrepancies in the camera modality.
Although the adoption of AE ensures low latency and clock synchronization, temporal misalignment can still occur when data is captured at different sensors at different time points (timestamps), especially due to their varying sampling rates. Some industry-grade software stacks (such as Autoware [^1] https://github.com/autowarefoundation/autoware and Apollo [^2] https://github.com/ApolloAuto/apollo) process the unimodal data independently and use post-processing techniques, such as Kalman filters, to ensure temporal alignment during the fusion based on their corresponding timestamps. Therefore, the efficacy of MMF-based perception relies on two fundamental assumptions. First, $A_{1}$: AE with TSN is secure, which ensures the sensor networks are clock-synchronized and have low latency. Second, $A_{2}$: the timestamp attached to the packet is trustworthy and can be used during the fusion process. However, in this work we propose that an attacker can exploit at least any of $A_{1}$ or $A_{2}$ to compromise the sensor network and expose the MMF to temporal misalignment.
MMF often relies on the assumption that integrating diverse sensor inputs will compensate for individual weaknesses and make it robust against any unimodal attack due to the redundancy of the collected data [8]. However, research shows this redundancy does not inherently provide superior robustness. Both unimodal and multimodal systems can be vulnerable to carefully designed unimodal adversarial or patch attacks [16, 17, 18, 19, 20, 21]. Moreover, research has shown that existing MMF-based perception models lack temporal robustness—the ability to produce reliable output despite temporal inconsistencies—against temporal misalignment and are quite fragile against a tiny delay, even in just one modality [10, 22, 23].
Therefore, to overcome the high overhead of the adversarial or patch attacks but still jeopardize the perception of MMF, we propose a network-based temporal misalignment attack named DejaVu. Under the DejaVu attack, an attacker compromises the in-vehicle network and maliciously creates a misaligned temporal data stream to cause an illusion during perception, even without crafting data. The attacker can achieve this by breaking any of the assumptions that the reliability of MMF is based upon. First, the attacker can break $A_{1}$ by exploiting the existing vulnerabilities in gPTP, and desynchronize sensor networks. PTP has been shown to be vulnerable to various cyberattacks, including man-in-the-middle (MITM) attacks [24], packet delay attacks [25], and spoofing [26, 27]. An attacker who gains such access to the in-vehicular network can manipulate PTP messages by introducing artificial delays, altering timestamps, or replaying old synchronization packets. This can cause different sensors to perceive capture data at slightly misaligned times, leading to misaligned sensor fusion.
Similarly, the attacker can break $A_{2}$ by exploiting the known vulnerabilities in network infrastructure to introduce intensional delays or distortions in the timestamps, causing temporal misalignment in the pre-processing phase of MMF. Particularly, automotive networks typically use UDP [13], a protocol inherently susceptible to man-in-the-middle (MITM) attacks, where attackers can intercept and manipulate communication flows and timestamps without accessing the encrypted data to exacerbate the desynchronization [28, 29]. In both cases, the attacker will be able to break the MMF assumptions of synchronized sensors or trustworthy timestamps, plummeting the perception accuracy. Fig. 1 illustrates how temporal delays due to DejaVu attacks in one of the modalities can degrade the performance of the MVXNet model [30]—a MMF-based 3D object detection model, potentially leading to unsafe driving conditions in AD.
The existing defense techniques against temporal misalignment assume benign timestamps and create alignment based on that. However, they will lack effectiveness against DejaVu, as the timestamps can be compromised as well. Hence, we propose a defense against DejaVu attack by extending this paradigm to include data content analysis along with the timestamps. In that regard, we propose Aion, a defense patch based on multimodal shared representation learning and dynamic time warping to detect and analyze the cross-modal temporal consistency. Aion works as a patch to any existing MMF models, and capitalizes on the semantic coherence of sensor data pairs, enabling the detection of inconsistencies that may arise from malicious temporal desynchronization. In summary, we make the following key contributions:
@@
  • We propose DejaVu, a novel temporal misalignment attack against MMF in AD that exploits vulnerabilities of in-vehicle networks and the fragility of multimodal fusion by selectively delaying sensor streams to disrupt perception in safety-critical tasks.
@@
  • We conduct a comprehensive empirical evaluation of DejaVu on state-of-the-art 3D object detection models (MVXNet [30], BEVFusion [31]) and a multi-object tracking model (MMF-JDT [32]) under various misalignment scenarios using the KITTI [33] and nuScenes [34] datasets. Our findings reveal distinct modality-specific vulnerabilities: object detectors are highly sensitive to LiDAR delays, while the tracking model is significantly impacted by camera timing disruptions. A single-frame LiDAR delay reduces 3D detection mAP by up to 88.5%, and a three-frame camera delay drops multiple object tracking accuracy (MOTA) by 73%.
  • To mitigate these attacks, we introduce Aion, a defense patch that leverages multimodal shared representation learning and dynamic time warping to detect temporal inconsistencies. Aion achieves AUROC scores of 0.92–0.98 with a low false positive rate, demonstrating robust and effective detection of temporal misalignment.
@@ II Related Work
II-A Attacks against MMF
The attacks against MMF can be classified into the following two types:
Sensor Spoofing Attacks.  Sensor spoofing attacks focus on feeding false data to a vehicle’s sensors to mislead the autonomous system. For instance, LiDAR sensors can be fooled by malicious laser signals that fabricate or erase objects in the sensed environment [18]. Beyond lasers, researchers have even created electromagnetic interference to tamper with LiDAR perception, achieving fake point insertions or deletions  [35]. Camera sensors are likewise vulnerable; strategically placed adversarial patches on road signs can cause a ‘Stop sign’ to be ignored or misclassified as another sign [36]. Such single-sensor spoofing attacks can trigger hazardous behaviors – e.g., a phantom obstacle can induce emergency braking, or a hidden obstacle can go unnoticed – underscoring how an attacker controlling one input modality can compromise the whole system.
Such unimodal sensor spoofing attacks can be launched against multimodal model, while the remaining modalities remain unattacked. For example, the attacker can inject fake LiDAR points to create fake objects, aligning them to the camera’s view frustum [19], so it remains consistent with the camera image to avoid inconsistencies during the sensor fusion. On the other hand, it was assumed that an adversarial attack against both sensors at once is unlikely to be successful against AD perception that has multimodal inputs such as camera and LiDAR data. However, researchers have devised attacks that break this assumption by crafting inputs that confound multiple modalities simultaneously. It is possible to generate an adversarial 3D physical object that appears innocuous to humans but fools the fused camera+LiDAR detection pipeline [37]. Attackers have even targeted high-level multi-modal representations by perturbing an image or audio input to make its fused embedding mimic an arbitrary target [38]. These sensor spoofing attacks demonstrate that simply having multiple sensors is not sufficient. With carefully constructed inputs, an adversary can simultaneously mislead camera and LiDAR sensors, defeating the very redundancy meant to ensure safety.
Temporal Delay Attack.  While time delay attacks have not been extensively studied within AD, they have created significant attention in other cyber-physical systems (CPS) domains, such as power systems, wireless networks, unmanned aerial vehicles (UAVs), robotics, and time-sensitive networks (TSNs). In power systems, the PTP is crucial for synchronization measurements and events. However, its vulnerability to time delay attacks poses risks to grid stability [39, 40]. Wireless networks, particularly those used in industrial control systems, are susceptible to time delay attacks that can disrupt communication and control processes [41]. In the realm of UAVs, time delay attacks can significantly impact flight control and mission execution [42, 43]. Moreover, multimodal temporal misalignment in sensor fusion has been shown to degrade the accuracy of simultaneous localization and mapping (SLAM) [44]. Software timing interference can also be exploited to cause system destabilization in CPS [45]. Time-delay attacks also impact the traffic transmission quality under different combinations of Time-Aware Shaper (TAS) or Cyclic Queue Forwarding (CQF) in TSN [46]. Furthermore, vulnerabilities in AE can allow for command injection attacks, leading to replay feed in camera [47]. However, none of the existing works consider the impact of temporal misalignment against task-agnostic MMF-based perception.
@@ II-A Attacks against MMF
II-B Defense Strategies
In the realm of autonomous vehicle security, various defense mechanisms have been proposed to counteract sensor spoofing and multimodal fusion attacks. These defenses can be broadly categorized into spatiotemporal consistency checks, specification-aware recovery strategies, and hardware-based techniques [48]. One prominent approach involves detecting inconsistencies across different sensor modalities over time. In this defense mechanism, spatiotemporal consistency is checked to identify anomalies in sensor data, such as discrepancies in object motion or position between LiDAR and camera inputs [49]. On the hardware side, manufacturers are improving sensor resilience: newer LiDAR models employ techniques such as pulse signal authentication or unique scanning patterns to make spoofed echoes easier to spot [20]. However, none of these provide the means to detect temporal misalignment attacks in multimodal fusion for AD.

### v2

# Temporal Misalignment Attacks against Multimodal Perception in Autonomous Driving

@@ Abstract
Multimodal fusion (MMF) plays a critical role in the perception of autonomous driving, which primarily fuses camera and LiDAR streams for a comprehensive and efficient scene understanding. However, its strict reliance on precise temporal synchronization exposes it to new vulnerabilities. In this paper, we introduce DejaVu, an attack that exploits the in-vehicular network and induces delays across sensor streams to create subtle temporal misalignments, severely degrading downstream MMF-based perception tasks. Our comprehensive attack analysis across different models and datasets reveals the sensors’ task-specific imbalanced sensitivities: object detection is overly dependent on LiDAR inputs, while object tracking is highly reliant on the camera inputs. Consequently, with a single-frame LiDAR delay, an attacker can reduce the car detection mAP by up to 88.5%, while with a three-frame camera delay, multiple object tracking accuracy (MOTA) for car drops by 73%. We further demonstrated two attack scenarios using an automotive Ethernet testbed for hardware-in-the-loop validation and the Autoware stack for end-to-end AD simulation, demonstrating the feasibility of the DejaVu attack and its severe impact, such as collisions and phantom braking.
@@ I Introduction
Autonomous driving (AD) is designed to navigate and interact with complex environments—a capability fundamentally reliant on a comprehensive understanding of its surroundings. The adoption of heterogeneous sensors, such as camera, LiDAR, and radar, that capture data from different modalities allows an accurate perception with enhanced accuracy and robustness [1]. Each individual modality has unique strengths; for example, cameras capture rich semantic details, LiDAR provides accurate depth measurements, and radar particularly excels in detecting speed, even in adverse weather conditions [2]. However, these sensors also face inherent limitations, such as a camera’s sensitivity to lighting variations, LiDAR’s lack of texture information, and radar’s sparsity, which can compromise performance when used independently [3, 4, 5]. Multimodal fusion (MMF)—the process of integrating multiple unimodal sensor data into a single and comprehensive representation—compensates for such individual sensor weaknesses, ensures accurate and robust perception, and efficient downstream tasks in AD [6, 7].
MMF remains a fundamental research challenge, primarily due to the heterogeneity across sensing modalities, including differences in data formats, spatial resolutions, and temporal characteristics [8, 9]. However, the robustness of temporal alignment—ensuring that all sensor data being fused corresponds to the same point in time—has received comparatively limited attention [10], despite being the fundamental enabler of MMF. Particularly, in AD, temporal alignment is a fundamental prerequisite to enable real-time perception, which directly informs decision-making and control. Misalignments, caused, for instance, by delays in specific sensor streams, can degrade perception performance, leading to object misdetection, localization errors, and scene misinterpretation. These inaccuracies can propagate downstream, affecting core functions such as control, maneuver planning, and safety interventions, ultimately compromising the reliability and safety of the vehicle [11]. Moreover, temporal alignment in MMF for AD remains a complex problem with the following challenges.
@@
Figure 1: Impact of DejaVu attack on 3D object detection. (a-b) show two benign scenarios without any temporal misalignment, and hence, accurate object detection. (c-d) Illustrate different temporal alignments between camera and LiDAR inputs and highlight how delayed sensor data can lead to incorrect detections, either by detecting non-existent objects (false positives) or missing present ones (false negatives). In (c), the MMF prioritizes the (delayed) LiDAR data and predicts three objects, including a pedestrian who is not present in the current camera view, resulting in seemingly accurate results from LiDAR’s perspective but a false detection from the camera’s perspective. However, in (d), the MMF still prioritizes the (updated) LiDAR data and predicts two objects, excluding the pedestrian who is present in the current camera view, resulting in a missed detection from the camera’s perspective. In both temporal misalignment attack cases, MMF biased its fusion toward LiDAR, failing to account for semantic discrepancies in the camera modality.
Clock synchronization.  To ensure temporally aligned perception in AD, sensor data from all the modalities must be timestamped relative to a common global clock with minimal drift. In practice, the electronic controller units (ECUs) hosting the sensors must be tightly synchronized—often at sub-microsecond precision required by AUTOSAR [12])—to enable accurate fusion and downstream reasoning. To achieve that, automotive Ethernet (AE) equipped with Time-Sensitive Networking (TSN) has become the de facto backbone of modern in-vehicle networks. TSN leverages the generalized Precision Time Protocol (gPTP, IEEE 802.1AS) to maintain global clock synchronization among distributed ECUs [13, 14]. Consequently, a core security assumption is: $\scriptstyle A_{1}$ The gPTP-based synchronization infrastructure is secure and maintains consistent global time across all ECUs. However, despite its widespread adoption in AE, gPTP was primarily designed to provide deterministic time synchronization—not to operate under adversarial conditions. Hence, gPTP lacks built-in security mechanisms, particularly cryptographic authentication, leaving it vulnerable to attacks such as grandmaster spoofing, delay injection, replay, and false time advertisement [15, 16, 17]. Such attacks can compromise the temporal alignment of sensor fusion, undermining the integrity of MMF-based perception, without any direct alteration to the sensor data or sensor ECU itself.
Timestamp Integrity.  Even when sensors are nominally synchronized, they often capture data at slightly different times due to variations in sampling rates caused by inherent hardware limitations [10]. Consequently, the timestamps attached to sensor messages become critical in multimodal fusion pipelines, as they serve as the primary reference for aligning asynchronous data streams by identifying the most temporally proximal pairs from buffered queues. This reliance gives rise to a fundamental security assumption: $\scriptstyle A_{2}$ Sensor ECUs are trusted entities that always provide accurate timestamps. However, this assumption can be violated. In vehicular systems, sensor ECUs may be compromised through remote exploits of unpatched software vulnerabilities [18, 19, 20], insecure Over-The-Air (OTA) updates [21, 22], or direct physical access via the OBD-II port [23, 24, 18]. Once compromised, an attacker-controlled ECU can inject messages containing legitimate sensor data but with forged timestamps. Since the fusion ECU uses timestamps as the basis for temporal alignment, this can result in the selection of data pairs that appear temporally aligned but are actually misaligned, thereby degrading the integrity of the sensor fusion process.
Middleware Integrity.  Sensor data exchange in production-grade AD (such as Autoware [^1] https://github.com/autowarefoundation/autoware) is often facilitated by robotic middleware frameworks, such as robot operating system (ROS [^2] https://github.com/ros2/ros2), where the data distribution service (DDS) serves as the underlying communication backbone. DDS enables a real-time and scalable publish-subscribe communication model for data sharing among distributed ECUs, making it a widely adopted solution for managing the high-bandwidth, low-latency communication demands of multimodal perception and control pipelines. Consequently, a third foundational assumption arises: $\scriptstyle A_{3}$ The ROS-based DDS infrastructure is secure and ensures the integrity, authenticity, and freshness of shared data. However, the design of ROS prioritizes performance and scalability over robust security guarantees. Since ROS does not enforce strong authentication, encryption, or source verification by default, ROS is vulnerable to a wide range of network-level and application-level attacks, including message spoofing, replay, and impersonation attacks [25]. As a result, an attacker with access to the ROS communication graph, for example, can impersonate legitimate nodes (e.g., a LiDAR publisher) and can publish fabricated sensor messages with targeted timestamps (i.e., replay attacks), which can be considered by the fusion ECU and mislead downstream perception tasks.
Moreover, prior work has demonstrated that existing MMF-based perception systems in AD lack temporal robustness—the ability to maintain reliable outputs in the presence of temporal inconsistencies—making them vulnerable even to benign, non-malicious misalignments [10, 26, 27]. In particular, perception pipelines have been shown to degrade significantly under small delays in just one modality. Existing studies, however, are limited in both scope and threat modeling: they primarily focus on random, system-induced delays and evaluate only a narrow subset of perception tasks—most commonly 3D object detection, without realizing them on any end-to-end AD software stacks. To investigate the true fragility of MMF-based perception under adversarial conditions, we present a comprehensive study of temporal misalignment attacks, which we term DejaVu. These attacks are realized by violating one or more of the core trust assumptions. Unlike prior work, we target both 3D object detection and multi-object tracking (MOT), and evaluate the effects of different delay distributions across multiple perception models and datasets. Fig. 1 illustrates how temporal delays due to DejaVu attacks in one of the modalities can degrade the performance of the MVXNet model [28]—a MMF-based 3D object detection model, potentially leading to unsafe driving conditions in AD.
In summary, we make the following key contributions:
@@
  • We propose DejaVu, a temporal misalignment attack against MMF in AD that exploits vulnerabilities of in-vehicle networks and the fragility of multimodal fusion by selectively delaying sensor streams to disrupt perception in safety-critical tasks.
@@
  • We conduct a comprehensive empirical evaluation of DejaVu on state-of-the-art 3D object detection models (MVXNet [28], BEVFusion [29]) using the KITTI [30] and nuScenes [31] datasets, and a multi-object tracking model (MMF-JDT [32]) under various misalignment scenarios using the KITTI [30] dataset. Our findings reveal distinct modality-specific vulnerabilities: object detectors are highly sensitive to LiDAR delays, while the tracking model is significantly impacted by camera timing disruptions. A single-frame LiDAR delay reduces 3D detection mAP by up to 88.5%, and a three-frame camera delay drops multiple object tracking accuracy (MOTA) by 73%.
  • To further validate our findings in a realistic autonomous driving setting, we built an automotive Ethernet testbed that models the sensor data acquisition and fusion pipeline. Using this platform, we implemented the DejaVu attack by violating $\scriptstyle A_{2}$, demonstrating its feasibility in a hardware-in-the-loop environment. Additionally, to assess the end-to-end consequences beyond perception—specifically on planning and control—we integrated DejaVu into Autoware, a production-grade, full-stack autonomous driving simulator, by breaking $\scriptstyle A_{3}$. In both environments, our experiments show that DejaVu is highly practical and can result in severe safety violations, including direct collisions and phantom braking events.
The remainder of the paper is organized as follows. Section II reviews related work; Section III outlines the threat model and proposes DejaVu attack; Section IV describes datasets and evaluation settings; Section V reports the evaluation results; Section VI reports the discussion and potential defenses; and Section VII concludes the paper.
@@ II Related Work
The fundamental research direction has been in the direction of spoofing sensors from a single modality, such as LiDAR [33, 34, 35] and camera [36] through different means of physical perturbation. Moreover, advanced attacks have demonstrated to even compromise multiple modalities together [37, 38, 39]. These sensor spoofing attacks demonstrate that simply having multiple sensors is not sufficient. With carefully constructed inputs, an adversary can simultaneously mislead camera and LiDAR sensors, defeating the very redundancy meant to ensure safety.
Unlike sensor spoofing attacks, time delay attacks have not been extensively studied within AD. They have created significant attention in other cyber-physical systems (CPS) domains, such as power systems [40, 41], wireless networks [42], unmanned aerial vehicles (UAVs) [43, 44, 15], and time-sensitive networks [45]. Software timing interference is also exploited to cause system destabilization in CPS [46]. Moreover, multimodal temporal misalignment in sensor fusion has been shown to degrade the accuracy of simultaneous localization and mapping (SLAM) [47], which was limited only to the fusion between IMU and camera data. Contrary to the existing works, we comprehensively study the impact of a temporal misalignment attack on task-agnostic MMF-based perception.
@@
In the realm of AD security, various defense mechanisms have been proposed to counteract sensor spoofing [48] and multimodal fusion attacks. These defenses can be broadly categorized into spatiotemporal consistency checks, specification-aware recovery strategies, and hardware-based techniques [49]. PercepGuard [50] detects misclassification attacks by enforcing consistency between object tracks and class labels, but it does not examine the temporal validity of sensor readings and thus cannot detect replayed LiDAR scans whose semantic trajectories remain plausible. Connecting the Dots [51] employs class‐specific autoencoders to uncover context violations introduced by adversarial perturbations, yet time‐shifted data aligns perfectly with learned scene co‐occurrence statistics and evades its checks. PhyScout [52] formalizes cross‐modal conflict detection to identify gross spoofing, but subtle timestamp manipulations within the synchronizer’s tolerance window introduce no overt spatial or modality discrepancies. These approaches overlook timestamp integrity and data freshness as a security property.